OpenStudy (anonymous):

how much fencing is needed to enclose a circular tulip garden whose area is 16pi sueare meters? 4pi m 16pi m 8pi m 32pi m

OpenStudy (amistre64):

area = pi r^2 16pi = pi r^2 16 = r^2 r = 4; now circumference = 2 pi r and thats what we want. 8pi m

OpenStudy (anonymous):

Curse you, beat me to it. I always thought wizards were good at math.

OpenStudy (amistre64):

you could always verify or unverify my results ;)

OpenStudy (anonymous):

okay thank you =)

OpenStudy (anonymous):

and haha wizards? na im bad at math, better with magic, my specialty =D

OpenStudy (anonymous):

Oh. I thought you were a mathemagician. I assumed you were gifted in the arts of Mathemagic.

OpenStudy (anonymous):

no sorry to dissapoint, my work deals more with things unkown to the human world...